# AgEval: A Benchmark for Zero-Shot and Few-Shot Plant Stress Phenotyping with Multimodal LLMs

### v1

@@ Abstract
Plant stress phenotyping traditionally relies on expert assessments and specialized models, limiting scalability in agriculture. Recent advances in multimodal large language models (LLMs) offer potential solutions to this challenge. We present AgEval, a benchmark comprising 12 diverse plant stress phenotyping tasks, to evaluate these models’ capabilities. Our study assesses zero-shot and few-shot in-context learning performance of state-of-the-art models, including Claude, GPT, Gemini, and LLaVA. Results show significant performance improvements with few-shot learning, with F1 scores increasing from 46.24% to 73.37% in 8-shot identification for the best-performing model. Few-shot examples from other classes in the dataset have negligible or negative impacts, although having the exact category example helps to increase performance by 15.38%. We also quantify the consistency of model performance across different classes within each task, finding that the coefficient of variance (CV) ranges from 26.02% to 58.03% across models, implying that subject matter expertise is needed – of ’difficult’ classes – to achieve reliability in performance. AgEval establishes baseline metrics for multimodal LLMs in agricultural applications, offering insights into their promise for enhancing plant stress phenotyping at scale. Benchmark and code can be accessed at:
https://anonymous.4open.science/r/AgEval/
@@ 1 Introduction
Food security stands as a paramount global challenge, necessitating sustainable enhancements in agricultural productivity to sustain a growing world population [1]. Agricultural research has increasingly turned to advanced technologies, particularly computer vision and artificial intelligence (AI), to optimize crop management to improve yield, profitability, and sustainability [2]. Phenotyping is the practice of visually inspecting plants and plant products to extract agronomically relevant features that help make downstream decisions. Traditional phenotyping methods—especially for plant stress phenotyping—involve expert visual inspections, are time-consuming, labor-intensive, often subjective, and fail to scale up. Computer vision can play a major role in plant stress phenotyping [3, 4].
@@
Most problems in the field of plant stress phenotyping can be divided into three primary categories: identification, classification, and quantification of plant stresses. In this study, we adopt an expansive view of plant stress phenotyping that includes traditional stresses, pest infestations, and seed quality issues, reflecting the interconnected nature of plant health factors in agricultural systems. Identification involves detecting the presence of stress—such as drought/heat stress, nutrient deficiency, pathogen infection, or pest/insect/weed infestation—in plants. Classification categorizes the stress into classes defined by subject matter experts, while quantification measures the severity or extent of the stress, including impacts on seed quality. Each task presents unique challenges and necessitates sophisticated analytical methods [5].
@@
Recent advances in computer vision and machine learning have created new opportunities for automating and enhancing plant stress phenotyping [6]. However, the development and deployment of AI-assisted phenotyping systems face several challenges. The effectiveness of these systems relies heavily on the quality and quantity of annotated data. Annotating plant stress data requires expert knowledge in plant pathology, entomology, agronomy, and other disciplines, making the process both costly and time-consuming. This creates a significant bottleneck in the development of AI models for agricultural applications.
@@
Researchers have explored various techniques to develop models that perform well with limited annotated data in plant stress phenotyping tasks [7]. These techniques include transfer learning from large-scale datasets such as ImageNet [8], self-supervised learning using contrastive methods on unlabeled plant images [7], and leveraging vision foundation models for improved generalization [9]. Despite these advancements, challenges persist due to the domain shift between general image datasets and agricultural imagery and the need for specialized visual understanding in plant stress phenotyping. The fine-grained nature of agricultural tasks often requires more nuanced feature recognition than what general pretrained models typically offer, underscoring the necessity for more adaptive and domain-specific approaches.
@@
Recent advances in multimodal large language models (LLMs) have shown promising results in image-text understanding tasks [10, 11]. These models demonstrate remarkable capabilities in processing and understanding visual and textual information jointly [12, 13, 14]. Additionally, these LLMs can perform few-shot in-context learning, adapting to new tasks with just a few examples in the prompt [15]. Their ability to leverage both visual and textual information could potentially bridge the gap between general visual understanding and domain-specific agricultural knowledge.
Our study aims to explore the potential of large multimodal vision-language models for plant stress phenotyping tasks. We hypothesize that their broad understanding of visual and textual information could be particularly beneficial in scenarios where annotated plant stress data is limited. By leveraging their pretrained knowledge of diverse datasets and their ability to understand natural language instructions, these models might offer a more flexible and adaptable solution for plant stress phenotyping and other agricultural AI tasks. In this work, we present a comprehensive evaluation of state-of-the-art multimodal LLMs on plant stress phenotyping tasks, focusing on the identification, classification, and quantification of plant stresses. We curate a benchmark dataset from open sources, reflecting real-world scenarios in agriculture. Our study aims to assess both zero-shot and few-shot in-context learning performance of these models on specialized agricultural tasks.
@@
This study makes several key contributions to plant stress phenotyping using multimodal Large Language Models (LLMs). First, it evaluates these models on plant stress phenotyping tasks, providing valuable insights into their potential to overcome the limitations of traditional approaches. Second, it introduces a curated benchmark dataset that closely mirrors real-world agricultural scenarios, offering a valuable resource for future research on adapting pretrained models to agricultural tasks. Third, the study analyzes the performance of few-shot in-context learning on specialized agricultural tasks, addressing the challenge of limited annotated data and assessing how effectively multimodal LLMs can leverage their pretrained knowledge. Fourth, it presents a comparative analysis of various state-of-the-art multimodal LLMs in the context of plant stress phenotyping. Finally, the study provides a quantitative assessment of how example relevance impacts few-shot in-context learning, further advancing our understanding of these models in agricultural applications.
@@ 2 Methodology
This section details our approach to evaluating multimodal LLMs for plant stress phenotyping tasks.
@@ 2.2 AgEval Benchmark Dataset Curation
We curated a diverse dataset comprising 12 subsets, each targeting specific plant stress phenotyping tasks. This collection, compiled from open-source resources, covers three main categories: Identification, Classification, and Quantification. These categories encompass various aspects of plant stress, from seed quality to pest infestations, reflecting the diverse challenges in agricultural stress assessment. The details of these datasets are provided in Table S1. We sampled 100 images from each dataset, evenly distributed across classes. The number of samples per class and the structure and relative sizes of the dataset components are illustrated in Figure S1.
Identification: We considered datasets addressing the identification of seed morphology variations due to stresses, foliar diseases, and pests including weeds and insects that cause stresses. The Durum Wheat Dataset [16] and Soybean Seeds dataset [17] support seed morphology tasks, which involve identifying stress-induced changes in seed characteristics. The Mango Leaf Disease Dataset [18, 19], Bean Leaf Lesions Classification dataset [20], and Soybean Diseases dataset [21] enable foliar stress tasks, focusing on identifying diverse plant stresses affecting leaves, including diseases and adverse environmental conditions. The DeepWeeds dataset [22] and Dangerous Farm Insects dataset [23] facilitate pest identification tasks, which involve recognizing weeds and insects that cause plant stress. These datasets and their associated tasks collectively contribute to assessing stress impacts on seed quality, disease management, and pest control strategies in agriculture.
@@ 2.3 Model Selection and Evaluation
We evaluated six multimodal LLMs: three state-of-the-art models (GPT-4V [34], Claude 3.5 Sonnet [35], and Gemini 1.5 Pro [36]), two budget-friendly options (Claude 3 Haiku [37] and Gemini 1.5 Flash [36]), and one open-source alternative (LLaVA v1.6 34B [38]). This selection encompasses a range of commercially available and open-source options to provide a comprehensive evaluation.
We evaluate LLM performance using both zero-shot and few-shot approaches. Zero-shot testing reveals inherent model biases, while few-shot testing (with 1, 2, 4, and 8 examples) indicates how model specialization might change with learning. For few-shot evaluations, we randomly select examples from the dataset. It’s important to note that our few-shot learning refers to in-context learning capability with a few examples, not fine-tuning in a few-shot manner [39].
@@ 2.4 Performance Metrics
$\text{NMAE}=\frac{\sum_{i=1}^{n}|y_{i}-\hat{y}_{i}|}{\max(y)-\min(y)}\cdot 100$ (2)
where $n$ is the number of samples, $y_{i}$ is the true value, $\hat{y}_{i}$ is the predicted value, and $\max(y)$ and $\min(y)$ are the maximum and minimum labels in the dataset. Note that the labels for classification problem are mapped to their corresponding ordinal values to be able to calculating NMAE. To handle out-of-vocabulary predictions from large language models in classification tasks, we assigned the worst possible score to unseen labels in the ordinal mapping. This approach penalizes predictions that fall outside the expected set of classes, which is important in agricultural applications where misclassification can lead to incorrect treatment decisions.
@@ 2.4.1 Relevance of Examples in Few-Shot Learning
This analysis (conducted for identification problems) quantifies the impact of example relevance in few-shot learning for the AgEval benchmark. We examine the effectiveness of examples from the same category versus diverse examples in predicting image labels. This analysis aims to determine whether examples from the same dataset (or in a real-world scenario, examples related to the farmer’s actual input) have a positive or negative impact on predictions. The findings aim to inform strategies for both end-users with limited examples and practitioners considering optimal few-shot prompts for multimodal language models. For this we utilized previous experiment runs where few shot examples and their labels were logged for each input across different shot settings.
@@
Table 1: 0-shot Performance of Large Language Models on AgEval Benchmark
@@ 2.4.2 Intra-task Uniformity
To evaluate the performance of LLM across different classes within individual (identification) tasks, we employ the Coefficient of Variation (CV) of all classes’ F1 scores. For example, a model could give high overall F1 score in identifying mangoe leaf disease but could be completely failing in identifying one of the class ’Cutting Weevil’. This analysis quantifies edge cases like that. For each model-dataset combination, we calculate the CV as follows:
$CV=\frac{\sigma}{\mu}\cdot 100$ (14)
where $\sigma$ is the standard deviation of F1 scores across classes, and $\mu$ is the mean F1 score. The CV provides a normalized measure of dispersion, allowing for comparison across datasets with different scales. For each identification dataset $d$ and model $m$, we calculate:
$CV_{d,m}=\frac{\sqrt{\frac{1}{n-1}\sum_{i=1}^{n}(F1_{i,d,m}-\overline{F1_{d,m}})^{2}}}{\overline{F1_{d,m}}}\cdot 100$ (15)
where $n$ is the number of classes in dataset $d$, $F1_{i,d,m}$ is the F1 score for class $i$, and $\overline{F1_{d,m}}$ is the mean F1 score across all classes for dataset $d$ and model $m$. We then compute average CVs across datasets and models:
@@
where $M$ is the set of all models and $D$ is the set of all datasets. These metrics allow us to quantify and compare the consistency of model performance across classes, identifying potential biases or challenges in specific classes or model behaviors.
@@ 2.5 Prompt Engineering
We designed task-specific prompts to guide the multimodal LLMs in performing the ICQ tasks. The prompts were structured to provide clear instructions and ensure consistent output formatting across all models.
Given the image, identify the class. Use the following list of possible classes for your prediction It should be one of the : {expected_classes}. Be attentive to subtle details as some classes may appear similar. Provide your answer in the following JSON format:{"prediction": "class_name"}Replace "class_name" with the appropriate class from the list above based on your analysis of the image. The labels should be entered exactly as they are in the list above i.e., {expected_classes}. The response should start with { and contain only a JSON object (as specified above) and no other text.
@@
Table 2: 8-shot Performance of Large Language Models on AgEval Benchmark
@@ 3.1 Zero-shot Performance
The zero-shot performance of various models on the AgEval benchmark reveals interesting patterns in their capabilities without task-specific training (Figure 2). In identification tasks, Gemini-pro-1.5 demonstrates the strongest performance with an MRR of 0.69, significantly outperforming other models. This suggests that Gemini-pro-1.5’s pretraining allows it to generalize well to AgEval’s identification tasks without additional context.
For classification and quantification tasks, GPT-4o emerges as the top performer with an MRR of 0.70, closely followed by Claude-3.5-sonnet at 0.54. This indicates that these models have a robust ability to understand and analyze images for more complex AgEval tasks like disease severity assessment and stress tolerance evaluation, even without specific examples. Interestingly, while Gemini-pro-1.5 excels in identification, it shows a notable drop in performance for classification and quantification tasks (MRR of 0.37). This disparity highlights the varying strengths of different models across task types in the AgEval benchmark, emphasizing the importance of model selection based on specific requirements within the agricultural domain.
Figure 2: Zero-shot comparative performance of LLMs.
Identification Tasks (F1 Score): In zero-shot identification tasks, Gemini-pro-1.5 leads with an average F1 score of 50.45 across the 7 tasks, outperforming others in 4 out of 7 tasks. It excels in Dangerous Insects (82.67) and Bean Leaf Lesions (77.22) identification. GPT-4o follows with 46.24 F1, showing strength in Mango Leaf Disease identification (58.41). Claude-3.5-sonnet performs competitively (46.02), particularly in Soybean Seeds identification (38.70). Notably, performance on the Soybean Diseases task is generally low across models, likely due to low image resolution.
Classification and Quantification Tasks (NMAE): For zero-shot classification and quantification, Claude-3.5-sonnet leads with the lowest average NMAE of 19.73 across 5 tasks, excelling in FUSARIUM 22 classification (18.25). GPT-4o follows closely (21.40), outperforming in 3 out of 5 tasks, including Yellow Rust 19 classification (17.19) and InsectCount quantification (15.80).

### v2

# Leveraging Vision Language Models for Specialized Agricultural Tasks

@@ Abstract
As Vision Language Models (VLMs) become increasingly accessible to farmers and agricultural experts, there is a growing need to evaluate their potential in specialized tasks. We present AgEval, a comprehensive benchmark for assessing VLMs’ capabilities in plant stress phenotyping, offering a solution to the challenge of limited annotated data in agriculture. Our study explores how general-purpose VLMs can be leveraged for domain-specific tasks with only a few annotated examples, providing insights into their behavior and adaptability. AgEval encompasses 12 diverse plant stress phenotyping tasks, evaluating zero-shot and few-shot in-context learning performance of state-of-the-art models including Claude, GPT, Gemini, and LLaVA. Our results demonstrate VLMs’ rapid adaptability to specialized tasks, with the best-performing model showing an increase in F1 scores from 46.24% to 73.37% in 8-shot identification. To quantify performance disparities across classes, we introduce metrics such as the coefficient of variation (CV), revealing that VLMs’ training impacts classes differently, with CV ranging from 26.02% to 58.03%. We also find that strategic example selection enhances model reliability, with exact category examples improving F1 scores by 15.38% on average. AgEval establishes a framework for assessing VLMs in agricultural applications, offering valuable benchmarks for future evaluations. Our findings suggest that VLMs, with minimal few-shot examples, show promise as a viable alternative to traditional specialized models in plant stress phenotyping, while also highlighting areas for further refinement. Results and benchmark details are available at: https://github.com/arbab-ml/AgEval
@@ 1 Introduction
Food security is a critical global challenge requiring sustainable improvements in agricultural productivity [1]. Agricultural research increasingly utilizes computer vision and AI to optimize crop management and enhance yield, profitability, and sustainability [2]. Phenotyping involves visually inspecting plants to extract agronomically relevant features. While traditional methods are time-consuming and labor-intensive, recent advancements in computer vision offer promising solutions to improve efficiency and scalability [3, 4]. These innovations present opportunities to revolutionize plant stress phenotyping, potentially leading to more objective, rapid, and large-scale assessments that could significantly boost agricultural productivity.
@@
Plant stress phenotyping tasks primarily fall into three categories: identification, classification, and quantification. Our study adopts a comprehensive view of plant stress phenotyping, encompassing traditional stresses, pest infestations, and seed quality issues. Identification detects stress presence (e.g., drought, nutrient deficiency, pathogens, pests). Classification categorizes stress into expert-defined classes, while quantification measures stress severity or extent, including seed quality impacts. Each task requires sophisticated analytical methods [5].
@@
Recent computer vision and machine learning advances offer new opportunities for automating plant stress phenotyping [6]. However, developing specialized models for agricultural applications faces challenges, primarily due to the need for high-quality annotated data. Expert knowledge in plant pathology, entomology, and agronomy is required for annotation, making it costly and time-consuming, thus creating a bottleneck in model development.
@@
Researchers have explored techniques to develop effective models with limited annotated data [7], including transfer learning from large-scale datasets [8], self-supervised learning on unlabeled plant images [7], and leveraging vision foundation models [9]. Despite progress, challenges remain due to domain differences between general and agricultural imagery, and the need for specialized visual understanding in plant stress phenotyping. The fine-grained nature of agricultural tasks often requires more nuanced feature recognition than general pretrained models offer, highlighting the need for adaptive, domain-specific approaches.
With traditional specialized model development, obtaining sufficient annotated data becomes increasingly challenging for each new use case. Typically, thousands of annotated examples are required to build effective models for specific agricultural tasks. As shown in Table S2, researchers have employed various specialized techniques to address this issue, including transfer learning, hybrid models, and custom architectures tailored to specific agricultural tasks. While these specialized models achieve impressive performance scores (e.g., 94% accuracy on Soybean Disease classification and 91% F1 score on DeepWeeds weed identification), they require significant domain expertise and computational resources to implement effectively. The high performance of these traditional approaches demonstrates their effectiveness for specific tasks, though their development and deployment costs remain substantial challenges.
@@
Recent advances in vision language models (VLMs) have shown promising results in image-text understanding tasks [10, 11]. These models demonstrate remarkable capabilities in processing and understanding visual and textual information jointly [12, 13, 14]. Additionally, VLMs can perform few-shot in-context learning (learning from a small number of examples without model updates), adapting to new tasks with just a few examples in the prompt [15]. VLMs often require only a couple of examples to learn and perform well on new tasks, potentially offering a more efficient solution to the data scarcity problem in agricultural applications.
Our study explores VLMs’ potential for plant stress phenotyping tasks, hypothesizing that their broad visual-textual understanding could benefit scenarios with limited annotated data. We present a comprehensive evaluation of state-of-the-art VLMs on identification, classification, and quantification of plant stresses, using a curated benchmark dataset reflecting real-world agricultural scenarios. As these models become increasingly accessible to agricultural stakeholders, our study assesses their reliability and effectiveness in specialized tasks, providing insights into their practical potential for plant stress phenotyping.
@@
This study makes several key contributions to plant stress phenotyping using VLMs. First, it evaluates these models on plant stress phenotyping tasks, providing insights into their potential to overcome traditional approach limitations. Second, it introduces a curated benchmark dataset mirroring real-world agricultural scenarios. Third, the study analyzes few-shot in-context learning performance on specialized agricultural tasks. Fourth, it presents a comparative analysis of various state-of-the-art VLMs in plant stress phenotyping. Finally, the study provides a quantitative assessment of how example relevance impacts few-shot in-context learning, advancing our understanding of these models in agricultural applications.
@@ 2 Methodology
This section details our approach to evaluating vision language models (VLMs) for plant stress phenotyping tasks.
@@ 2.2 AgEval Benchmark Dataset Curation
We curated a diverse dataset comprising 12 subsets, each targeting specific plant stress phenotyping tasks. This collection, compiled from open-source resources, covers three main categories: Identification, Classification, and Quantification. These categories encompass various aspects of plant stress, from seed quality to pest infestations, reflecting the diverse challenges in agricultural stress assessment. The details of these datasets are provided in Table S1. We sampled 100 images from each dataset, evenly distributed across classes. Further details provided in Figure S1.
Identification: We considered datasets addressing the identification of seed morphology variations due to stresses, foliar diseases, and pests including weeds and insects that cause stresses. The Durum Wheat Dataset [16] and Soybean Seeds dataset [17] support seed morphology tasks, which involve identifying stress-induced changes in seed characteristics. The Mango Leaf Disease Dataset [18, 19], Bean Leaf Lesions Classification dataset [20], and Soybean Diseases dataset [21] enable foliar stress tasks, focusing on identifying diverse plant stresses affecting leaves, including diseases and adverse environmental conditions. The DeepWeeds dataset [22] and Dangerous Farm Insects dataset [23] facilitate pest identification tasks, which involve recognizing weeds and insects [24] that cause plant stress. These datasets and their associated tasks collectively contribute to assessing stress impacts on seed quality, disease management, and pest control strategies in agriculture.
@@ 2.3 Model Selection and Evaluation
We evaluated six vision language models (VLMs): three state-of-the-art models (GPT-4o [35], Claude 3.5 Sonnet [36], and Gemini 1.5 Pro [37]), two budget-friendly options (Claude 3 Haiku [38] and Gemini 1.5 Flash [37]), and one open-source alternative (LLaVA v1.6 34B [39]). This selection encompasses a range of commercially available and open-source options to provide a comprehensive evaluation.
We evaluate VLM performance using both zero-shot and few-shot approaches. Zero-shot testing reveals inherent model capabilities, while few-shot testing (with 1, 2, 4, and 8 examples) demonstrates the models’ ability to adapt to new tasks with minimal examples. For few-shot evaluations, we randomly select examples from the dataset. It’s important to note that our few-shot learning refers to in-context learning capability with a few examples, not fine-tuning in a few-shot manner [40].
@@ 2.4 Performance Metrics
$\text{NMAE}=\frac{\sum_{i=1}^{n}|y_{i}-\hat{y}_{i}|}{\max(y)-\min(y)}\cdot 100\%$ (2)
where $n$ is the number of samples, $y_{i}$ is the true value, $\hat{y}_{i}$ is the predicted value, and $\max(y)$ and $\min(y)$ are the maximum and minimum labels in the dataset. Note that the labels for classification problem are mapped to their corresponding ordinal values to be able to calculating NMAE. To handle out-of-vocabulary predictions from vision language models in classification tasks, we assigned the worst possible score to unseen labels in the ordinal mapping. This approach ensures consistency in evaluating predictions across all models.
@@ 2.4.1 Relevance of Examples in Few-Shot Learning
This analysis explores the impact of example relevance in few-shot learning for identification tasks. We investigate how examples from the same category versus diverse examples affect vision language models’ (VLMs) performance in predicting image labels. This study aims to understand whether examples from the same dataset (or in a real-world scenario, examples related to the farmer’s actual input) positively influence predictions, and whether related information steers the model towards more accurate category-specific predictions.
Our analysis utilizes data from previous experiment runs, where few-shot examples and their labels were logged for each input across different shot settings.
@@
Intuitively, a bullseye shot occurs when at least one of the k example images provided in the prompt belongs to the same class as the target image being evaluated.
Note: A bullseye shot requires at least one example to match the true label, not necessarily all k examples.
This definition helps us quantify the effectiveness of providing relevant examples, offering insights into how VLMs can rapidly adapt to specialized agricultural tasks with minimal context, and how related information might enhance prediction accuracy for specific categories.
@@
Table 1: 0-shot Performance of VLMs on AgEval Benchmark, Models Sorted by Average Performance (Highest to Lowest)
@@ 2.4.2 Intra-task Uniformity
To evaluate the performance consistency of Vision Language Models (VLMs) across different classes within individual (identification) tasks, we employ the Coefficient of Variation (CV) of all classes’ F1 scores. This analysis is crucial for understanding the robustness and reliability of VLMs in domain-specific applications, not just in agriculture but across various specialized fields. By quantifying performance variability, we can identify potential biases or gaps in model training that may impact real-world deployment.
For each model-dataset combination, we calculate the CV as follows:
$CV=\frac{\sigma}{\mu}\cdot 100\%$ (14)
where $\sigma$ is the standard deviation of F1 scores across classes, and $\mu$ is the mean F1 score. The CV provides a normalized measure of dispersion, allowing for comparison across datasets with different scales.
For each identification dataset $d$ and model $m$, we calculate:
$CV_{d,m}=\frac{\sqrt{\frac{1}{n-1}\sum_{i=1}^{n}(F1_{i,d,m}-\overline{F1_{d,m}})^{2}}}{\overline{F1_{d,m}}}\cdot 100\%$ (15)
where $n$ is the number of classes in dataset $d$, $F1_{i,d,m}$ is the F1 score for class $i$, and $\overline{F1_{d,m}}$ is the mean F1 score across all classes for dataset $d$ and model $m$.
We then compute average CVs across datasets and models:
@@
where $M$ is the set of all models and $D$ is the set of all datasets.
These metrics allow us to quantify and compare the consistency of model performance across classes, highlighting areas where VLMs excel and identifying potential challenges in specific classes or model behaviors. This analysis provides valuable insights into the adaptability and robustness of VLMs in handling diverse tasks, which is essential for their effective application in specialized domains beyond agriculture, such as medical imaging, industrial quality control, or environmental monitoring.
@@ 2.5 Prompt Engineering
We designed task-specific prompts to guide the VLMs in performing the ICQ tasks. The prompts were structured to provide clear instructions and ensure consistent output formatting across all models.
(keyval) Package keyval Error: breaklines undefinedSee the keyval package documentation for explanation.I can’t help(keyval) Package keyval Error: breakanywhere undefinedSee the keyval package documentation for explanation.I can’t help
Given the image, identify the class. Use the following list of possible classes for your prediction It should be one of the : {expected_classes}. Be attentive to subtle details as some classes may appear similar. Provide your answer in the following JSON format:{"prediction": "class_name"}...
@@
Table 2: 8-shot Performance of VLMs on AgEval Benchmark, Models Sorted by Average Performance (Highest to Lowest)
@@ 3.1 Zero-shot Performance
The zero-shot performance of Vision Language Models (VLMs) on the AgEval benchmark reveals interesting patterns (Figure 2). In identification tasks, Gemini-pro-1.5 demonstrates the strongest performance with an MRR of 0.69. For classification and quantification tasks, GPT-4o emerges as the top performer with an MRR of 0.70, closely followed by Claude-3.5-sonnet at 0.54. This highlights the varying strengths of different models across task types, emphasizing the importance of model selection based on specific requirements within the agricultural domain.
The strong performance of larger models like GPT-4o and Gemini-pro-1.5 suggests that general-purpose training translates well to domain-specific performance, underscoring the value of comprehensive pretraining in enabling VLMs to adapt to specialized agricultural tasks without additional fine-tuning.
Identification Tasks (F1 Score): In zero-shot identification tasks, Gemini-pro-1.5 leads with an average F1 score of 50.45 across the 7 tasks, outperforming others in 4 out of 7 tasks. GPT-4o follows with 46.24 F1, showing strength in Mango Leaf Disease identification (58.41). Claude-3.5-sonnet performs competitively (46.02), particularly in Soybean Seeds identification (38.70).
Figure 2: Zero-shot comparative performance of VLMs.
Figure 3: 8-shot comparative performance of VLMs.
Classification and Quantification Tasks (NMAE): For zero-shot classification and quantification, Claude-3.5-sonnet leads with the lowest average NMAE of 19.73 across 5 tasks. GPT-4o follows closely (21.40), outperforming in 3 out of 5 tasks. These results reinforce that larger VLMs can effectively leverage their general-purpose training to perform well on specialized agricultural tasks, even without domain-specific fine-tuning.
Figure 5: Performance comparison on individual tasks of the AgEval benchmark across different shot settings (0 to 8 shots) for top-4 performing VLMs.